OpenStudy (anonymous):

given the equation: x^4+y^4=16. find dy/dx

OpenStudy (anonymous):

First you need to solve in order to Y. y^4 = 16-x^4 y= sqrt(sqrt(16-x^4)) now you just need to derivate: dy/dx = -x^3/(16-x^4)^(3/4)

OpenStudy (anonymous):

what is the d^2y/dx^2

OpenStudy (anonymous):

-(48 x^2)/(16-x^4)^(7/4)

OpenStudy (anonymous):

how did you get that?

OpenStudy (anonymous):

try this \[4x^3+4y^3y'=0\] \[y'=-\frac{x^3}{y^3}\]

OpenStudy (anonymous):

don't need to solve for y if you do not mind having a y in your answer

OpenStudy (anonymous):

That is correct.

OpenStudy (anonymous):

dy/dx(x^4)+dy/dx(y^4)=dy/dx(16) 4x^3+4y^3.dy/dx=0 dy/dx=-4x^3/4y^3 dy/dx=-x^3/y^3